# Rotational Properties of Inverted Hybrid Stars

### v1

@@ Abstract
We study the rotational properties of inverted hybrid stars (also termed cross stars), which have been recently proposed as a possible new class of compact stars characterized by an outer layer of quark matter and a core of hadrons, in an inverted structure compared to traditional hybrid stars. We analyze distinct models representing varying depths of quark-hadron phase transitions. Our findings reveal that, while cross stars rotating at their Kepler frequencies typically exhibit a significantly higher mass and larger circumferential radius as anticipated, interestingly, there is a significant increase in potential twin configurations in the case of rapid rotations. We further study sequences of constant baryonic mass, representing potential paths of rotational evolution. Our results indicate that not all stars in these sequences are viable due to the onset of phase transitions during spin-down, leading to possible mini-collapses. We also investigate the phenomenon of “back-bending” during spin-down sequences, which is manifested in a rather different shape for cross stars due to their inverted structure and the large density discontinuity caused by the strong phase transition, in contrast to traditional hybrid stars. Our research enriches existing studies by introducing the significant aspect of rotation, unveiling intriguing phenomena that could serve as observational markers.
@@ I Introduction
The concept of phase transitions within compact stars is a very realistic possibility that has been seriously considered for decades [1, 2, 3, 4, 5, 6, 7, 8]. Typically, based on our understanding of the baryonic matter phase diagram, the transition from hadronic matter (HM) to deconfined quark matter (QM) at some point within the compact object is considered, with the precise point of this transition is contingent upon the equation of state (EOS) of both the hadronic and quark phases [9, 10, 11, 12, 13, 14, 15]. A significant challenge in the study of hybrid stars is determining the hadronic-quark matter transition at the low temperatures typical of compact stars. Furthermore, it remains to be conclusively determined whether the transition occurs at the densities found within the interiors of neutron stars, although it has been suggested that this is probable [16]. Compact stars that undergo a phase transition to quark matter at inner regions are commonly referred to as hybrid stars.
@@
Recently, the provocative possibility of “inverted” hybrid stars was proposed and examined [17, 18]. These objects are composed of an outer layer of quark matter, with the inner stellar region composed of ordinary hadronic matter, in opposite to what one expects for traditional hybrid stars, hence the name ”inverted” hybrid stars or cross stars (CrSs) considering that the inverted phase transition occurs when the chemical potentials of the two matter phases cross. This concept is supported by the hypothesis of absolutely stable QM at low pressure, as demonstrated by [19] for strange quark matter (SQM) that is composed of a nearly equal number of $u,d,s$ quarks, and [20] for up-down quark matter ($ud$QM) that composed of $u,d$ quarks, together with the possibility that at high pressures hadronic matter may be more stable than QM due to a smaller chemical potential in a large physical EOS parameter space [17]. This inverted phase transition cannot be excluded since it resides in the intermediate density and low-temperature regime where the QCD dynamics is non-perturbative and lattice calculation suffers from the fermion sign problem. Besides, as [17] observed, composition with $ud$QM allows a larger parameter space than those with SQM. As SQM can form strange quark stars, $ud$QM can also form up-down quark stars [21, 22, 23, 24, 25, 26, 27]. CrSs can be seen as configurations in which a hadronic core developed inside these quark stars.
@@
The properties of non-rotating (spherically symmetric) CrSs have been thoroughly examined in [17], with a focus on their macroscopic features such as mass and radius. This study also explored key quantities for potential merger scenarios, including tidal deformability. Later, the asteroseismology of CrSs was systematically studied in [18]. Building on the findings of [17], our current research delves deeper into the attributes of these objects, with a particular emphasis on the properties of rotating cross stars. The characteristics of rotating compact stars have garnered significant interest recently, particularly due to the potential emergence of high-angular momentum compact stars from merger events. We argue that an in-depth study of rotating CrSs is justified, as it could yield valuable insights for future dynamic simulations of post-merger remnants, in addition to its inherent scholarly value.
@@
This paper is organized as follows: Section I introduces the context and purpose of this work; Section II provides an overview of the microscopic model used in this study; Section III discusses our findings on the structure of rotating cross stars; Section IV explores potential evolutionary sequences of rotating stars; Section V describes the back-bending phenomena as observed in rotating cross stars; and Section VI offers our conclusions and future perspectives. In this paper, we adopt units where $G=c=1$.
@@ II Microscopic Model
In this study, we adopt the methodology of [17] and explore the concept of CrSs, which are theorized to be composed of quark matter at lower densities and hadronic matter at higher densities. As discussed in the introduction, we focus on the CrSs with $ud$QM, which has the grand potential expressed as follows:
@@
where $\mu$ is the average quark chemical potential. This can be understood as the grand potential of a free quark gas with QCD corrections parameterized by the parameter $a_{4}$, with $a_{4}=1$ signifying no corrections and larger values of $a_{4}$ denoting larger corrections [28]. $B$ is the effective bag constant. The parameter $\xi_{4}$ for $ud$QM, is given by [17] $\xi_{4}=\left[(1/3)^{4/3}+(2/3)^{4/3}\right]^{-3}\approx 1.86$.
From the potential given in Eq. (1) we can write the following equations for the energy per baryon of $ud$QM [25]
@@
while the equation of state (EOS) is given by
@@
with $\epsilon$ denoting the energy density.
The $ud$QM parameters must be defined within the $ud$QM hypothesis. As discussed in [17] we must take into account the absolute stability condition of QM at zero pressure, which constrains $(B,a_{4})$ via $(E/A)_{q}<930$ MeV, the energy per baryon number for most stable nucleus 56Fe. Therefore, following in the footsteps of [17], we consider the following benchmark values for the bag constant $B=20,35,50$ MeV/fm3 with the corresponding values of $a_{4}=0.35,0.62,0.88$, respectively. Each pair of values for parameters $B$ and $a_{4}$ will be denoted, heretofore as parameters set A, B, C. Explicitly, set A corresponds to $B=20$ MeV/fm3 and $a_{4}=0.35$, set B to $B=35$ MeV/fm3 and $a_{4}=0.62$, and set C to $B=50$ MeV/fm3 and $a_{4}=0.88$.
In the study of CrSs, it is necessary to consider a hadronic EOS. The authors of [17, 18] performed a comprehensive examination of a representative set of hadronic equations of states, which revealed similar qualitative features and small quantitative differences (for details, see their appendices). Consequently, following their observations, in this work we consider the APR EOS [29] for the hadronic sector, as sourced from the CompOSE database [30]. Employing the $ud$QM hypothesis, as well as the aforementioned hadronic EOS, quark matter transits to hadronic matter as pressure increases beyond the transition pressure $P_{\rm cr}$ as determined by the crossing of the chemical potentials of the two matter phases. The combined EOS for each benchmark model of CrSs is shown in Fig. 1.
Figure 1: The EOS for cross stars involves $ud$QM transitioning to hadronic matter (APR model) at high pressure. Parameter sets A, B, and C represent varying values for the $ud$QM bag constant and the $a_{4}$ parameter. Specifically, set A corresponds to $B=20$ MeV/fm3 and $a_{4}=0.35$, set B to $B=35$ MeV/fm3 and $a_{4}=0.62$, and set C to $B=50$ MeV/fm3 and $a_{4}=0.88$.
@@ III Structure of Rotating stars
The structure of rotating neutron stars presents a significantly more complex problem to solve compared to their non-rotating counterparts. For non-rotating, spherically symmetric neutron stars, the combination of Einstein’s equation and the equilibrium condition yields the renowned Tolman-Oppenheimer-Volkoff Equations (TOV) [31, 32].
@@
These equations must then be solved in all space, with flat space at infinity as boundary conditions. In this work we employ the code Astreus, which has been widely used to study rotating neutron stars [33, 34, 35, 36, 37]. This code employs the numerical method described in [38] and have been thoroughly explored in [39].
We show in Fig. 2 the mass as a function of the central density corresponding to the three different EOSs for CrSs that we have considered.
Figure 2: Mass as a function of central density is depicted for three different Equations of State (EOS) of compact stars (CrSs) studied. For each EOS, two curves are presented, represented by solid and dashed lines, corresponding to the spherical sequence where stars have no rotation, and the Kepler sequence where each star rotates at its maximum (Kepler) frequency, respectively.
@@
The findings illustrated in Fig. 2 reveal that CrSs undergoing a transition from quark matter to hadronic matter at lower densities tend to exhibit smaller masses. Conversely, transitions occurring at higher densities result in larger masses.
@@
It is also worth noting the emergence of twin star configurations, especially for CrSs where the transition occurs at lower densities. Twin star configurations refer to scenarios where two stable stars from the same family have identical mass but differ in central densities (and radii), resulting in varying compactness [40, 41, 42]. Typically, twin star setups involve the star with the smaller central density (less compact) being purely hadronic, while its more compact counterpart is hybrid. However, the case of inverted hybrid stars has the reversed situation, where the less compact twin is composed solely of quark matter, and the more compact twin contains both quark matter and hadronic matter, as explicitly shown in [17] for non-rotating CrSs. Here our new finding reveals that, when rotation is introduced, an intriguing phenomenon emerges: the twin branch significantly expands. This is particularly noteworthy in the case of EOS B, where a relatively pronounced twin branch appears when rotation is factored in. This phenomenon may also appear in traditional hybrid stars. For instance, Ref. [43] showed that adding rotations can introduce twin stars from normal stars.
The emergence of twin-star configurations becomes more pronounced when examining the Mass-Radius diagram for these CrS sequences. These findings are depicted in Figures 3, which display the gravitational mass as a function of the circumferential radius. The cyan-shaded region denotes the twin-star configurations. The insets in these figures illustrate the twin configurations for the corresponding non-rotating ones of the same EOS.
Figure 3: The mass as a function of circumferential radius for CrSs of EOS B (top panel) and EOS C (bottom panel) rotating at their Kepler frequencies. The blue shaded area denotes the zone where twin configurations occur. The insets illustrate the twin configurations of corresponding spherically symmetric (non-rotating) stars.
@@ IV Evolution Sequences
Given that a compact star may be born with a specific angular momentum, whether after a supernova or a merger event, exploring potential rotational evolution paths as they decelerate is warranted. This involves examining the evolution of a sequence with constant baryon mass from a state of high rotation to the non-rotating limit. Such sequences can reveal the potential evolutionary trajectories of isolated stars that decelerate without mass loss or gain, thereby preserving a constant baryonic mass.
@@
Figure 4: Gravitational mass as a function of circumferential radius for CrSs of EOS B. Each dashed curve represents evolution paths for stars with the indicated baryon mass. The gray-shaded region depicts the forbidden region above the Kepler sequence.
It is crucial to note that while the stars depicted by the constant baryonic sequences in Fig. 4 are hydrostatically stable, not all may be stable against oscillations. The stability condition for a sequence of stars with constant baryonic mass follows the turning point theorem (see for instance [44]) and can be written as
@@
where $J$ is the angular momentum and $\epsilon_{c}$ the central energy density. Besides the aforementioned condition, it is also necessary to consider the energy density gap in the EOS, which is evident in Fig. 1, resulting from the phase transition.
@@
Figure 6: Gravitational mass as a function of central density for the sequence of stars of EOS B. Orange shaded regions represent forbidden configurations.
As observed in Fig. 6, there exists a small region of unstable configuration near to the Kepler sequence. This suggests that stars may originate at lower frequencies (beyond the orange-shaded regions). Alternatively, if they are born with frequencies close to the Kepler limit, they will experience a sudden transition through the prohibited region. A more detailed analysis of the constant baryonic mass curves can provide a better understanding of this phenomenon, which can be facilitated with the aid of Fig. 7
Figure 7: Similar to Fig. 6, but with a zoomed-in region for each of the constant baryonic mass sequences. Points A and B in each figure denote the last stable configuration before and after the forbidden region.
@@
A similar result was outlined in reference [44]. One must bear in mind, however, that not all of this energy is to be radiated away, as part of the energy will remain in the star in the form of heat. Even if only 1% of the energy is radiated it could give rise to very energetic phenomena such as gamma-ray bursts and fast-radio bursts [45], thus warranting further investigations on the dynamics of such phenomena.

### v2

@@ Abstract
We study the rotational properties of inverted hybrid stars (IHSs), which have been recently proposed as a possible new class of compact stars characterized by an outer layer of quark matter and a core of hadronic matter, in an inverted structure compared to traditional hybrid stars. We analyze distinct models representing varying depths of quark-hadron phase transitions. Our findings reveal that, while IHSs rotating at their Kepler frequencies typically exhibit a significantly higher mass and larger circumferential radius as anticipated, interestingly, there is a significant increase in potential twin configurations in the case of rapid rotations. We further study sequences of constant baryonic mass, representing potential paths of rotational evolution. Our results indicate that not all stars in these sequences are viable due to the onset of phase transitions during spin-down, leading to possible mini-collapses. We also investigate the phenomenon of “back-bending” during spin-down sequences, which is manifested in a rather distinct shape for IHSs due to their inverted structure and the large density discontinuity caused by the strong phase transition. Our study extends to the braking index of IHSs, which could serve as observational hints for the detection of such objects. Furthermore, we also consider sequences of IHSs with constant angular momentum and demonstrate that they exhibit, with relatively high confidence, an approximate universal relation between supramassive configurations and the spherical mass limit. Furthermore, we also show that $\hat{I}-\hat{Q}$ relations found in previous studies also hold approximately for IHSs to some extent. Our research enriches existing studies by introducing the significant aspect of rotation, unveiling intriguing phenomena that could serve as observational markers.
@@ I Introduction
The concept of phase transitions within compact stars is a very realistic possibility that has been seriously considered for decades Dexheimer et al. (2018); Orsaria et al. (2019); Heiselberg and Hjorth-Jensen (1999); Glendenning (2001); Yang and Shen (2008); Reddy et al. (2000); Glendenning (1992); Burgio et al. (2002). Typically, based on our understanding of the baryonic matter phase diagram, the transition from hadronic matter (HM) to deconfined quark matter (QM) at some point within the compact object is considered, with the precise point of this transition is contingent upon the equation of state (EOS) of both the hadronic and quark phases Alford et al. (2005a); Zhang et al. (2023); Cierniak and Blaschke (2021); Peng et al. (2008); Blaschke et al. (2010); Dexheimer et al. (2012, 2013). A significant challenge in the study of hybrid stars is determining the hadronic-quark matter transition at the low temperatures typical of compact stars. Furthermore, it remains to be conclusively determined whether the transition occurs at the densities found within the interiors of neutron stars, although it has been suggested that this is probable Annala et al. (2020). Compact stars that undergo a phase transition to quark matter at inner regions are commonly referred to as hybrid stars.
@@
Recently, the provocative possibility of “inverted” hybrid stars was proposed and examined Zhang and Ren (2023); Zhang et al. (2024). These objects are composed of an outer layer of quark matter, with the inner stellar region composed of ordinary hadronic matter, in opposite to what one expects for traditional hybrid stars, hence the name “inverted” hybrid stars (IHSs) or cross stars by other names. This concept is supported by the hypothesis of absolutely stable QM at low pressure, as demonstrated by Witten (1984) for strange quark matter (SQM) that is composed of a nearly equal number of $u,d,s$ quarks, and Holdom et al. (2018) for up-down quark matter ($ud$QM) that composed of $u,d$ quarks [^1] Following Holdom et al. (2018), the properties of $ud$QM are in active explorations Iida and Fujie (2020); Xia et al. (2020, 2022); Bai and Korwar (2024); Bai and Chen (2024); Brown et al. (2024)., together with the possibility that at high pressures hadronic matter may be more stable than QM due to a smaller chemical potential in a large physical EOS parameter space Zhang and Ren (2023). This inverted phase transition cannot be excluded since it resides in the intermediate density and low-temperature regime where the QCD dynamics is non-perturbative and lattice calculation suffers from the fermion sign problem. Besides, as Zhang and Ren (2023) observed, composition with $ud$QM allows a larger parameter space than those with SQM. As SQM can form strange quark stars Alcock et al. (1986); Weber (2005); Zhang and Mann (2021); Pretel and Zhang (2024); Zhou et al. (2024), $ud$QM can also form up-down quark stars Zhang (2020); Ren and Zhang (2020); Cao et al. (2022); Wang et al. (2021); Li et al. (2022); Yuan et al. (2022). IHSs can be seen as configurations in which a hadronic core developed inside these quark stars.
@@
The properties of non-rotating (spherically symmetric) IHSs have been thoroughly examined in Zhang and Ren (2023), with a focus on their macroscopic features such as mass and radius. This study also explored key quantities for potential merger scenarios, including tidal deformability. Later, the asteroseismology of IHSs was systematically studied in Zhang et al. (2024). Building on the findings of Zhang and Ren (2023), our current research delves deeper into the attributes of these objects, with a particular emphasis on the properties of rotating IHSs. The characteristics of rotating compact stars have garnered significant interest recently, particularly due to the potential emergence of high-angular momentum compact stars from merger events. We argue that an in-depth study of rotating IHSs is justified, as it could yield valuable insights for future dynamic simulations of post-merger remnants, in addition to its inherent scholarly value.
@@
This paper is organized as follows: Section I introduces the context and purpose of this work; Section II provides an overview of the microscopic model used in this study; Section III discusses our findings on the structure of rotating IHSs; Section IV explores potential evolutionary sequences of rotating stars; Section V describes the back-bending phenomena as observed in rotating cross stars; and Section VI offers our conclusions and future perspectives. In this paper, we adopt units where $G=c=1$.
@@ II Microscopic Model
In this study, we adopt the methodology of Zhang and Ren (2023) and explore the concept of IHSs, which are theorized to be composed of quark matter at lower densities and hadronic matter at higher densities. As discussed in the introduction, we focus on the IHSs with $ud$QM, which has the grand potential expressed as follows:
@@
where $\mu$ is the average quark chemical potential. This can be understood as the grand potential of a free quark gas with QCD corrections parameterized by the parameter $a_{4}$, with $a_{4}=1$ signifying no corrections and smaller values of $a_{4}$ denoting larger corrections Alford et al. (2005b). $B$ is the effective bag constant. The parameter $\xi_{4}$ for $ud$QM, is given by Zhang and Ren (2023) $\xi_{4}=\left[(1/3)^{4/3}+(2/3)^{4/3}\right]^{-3}\approx 1.86$.
From the potential given in Eq. (1) we can write the following equations for the energy per baryon of $ud$QM Zhang and Mann (2021)
@@
and the chemical potential
$\mu_{\rm Q}=\frac{3\sqrt{2\pi}}{(a_{4}\xi_{4})^{1/4}}(P+B)^{1/4},$ (3)
while the equation of state is given by
@@
with $\epsilon$ denoting the energy density. Note that while the EOS for $ud$QM is independent of $a_{4}$, it will affect the chemical potential $\mu_{Q}$ and thus the point of quark-hadron phase transition.
The $ud$QM parameters must be defined within the $ud$QM hypothesis. As discussed in Zhang and Ren (2023) we must take into account the absolute stability condition of QM at zero pressure, which constrains $(B,a_{4})$ via $(E/A)_{q}<930$ MeV, the energy per baryon number for 56Fe, known to be the nucleus with the lowest mass per nucleon. Therefore, following in the footsteps of Zhang and Ren (2023), we consider the following benchmark values for the bag constant $B=20,35,50$ MeV/fm3 with the corresponding values of $a_{4}=0.35,0.62,0.88$, respectively. Each pair of values for parameters $B$ and $a_{4}$ will be denoted, heretofore as parameters set A, B, C. Explicitly, set A corresponds to $B=20$ MeV/fm3 and $a_{4}=0.35$, set B to $B=35$ MeV/fm3 and $a_{4}=0.62$, and set C to $B=50$ MeV/fm3 and $a_{4}=0.88$.
In the study of IHSs, it is necessary to consider a hadronic EOS. The authors of Zhang and Ren (2023); Zhang et al. (2024) performed a comprehensive examination of a representative set of hadronic equations of states, which revealed similar qualitative features and small quantitative differences (for details, see their appendices). Consequently, following their observations, in this work we consider the APR EOS a. Akmal et al. (1998) for the hadronic sector, as sourced from the CompOSE database COMPOSE . Employing the $ud$QM hypothesis, as well as the aforementioned hadronic EOS, quark matter transits to hadronic matter as pressure increases beyond the transition pressure $P_{\rm cr}$ as determined by the crossing of the chemical potentials of the two matter phases. The combined EOS for each benchmark model of IHSs is shown in Fig. 1.
Figure 1: The EOS for inverted hybrid stars involves $ud$QM transitioning to hadronic matter (APR model) at high pressure. Parameter sets A, B, and C represent varying values for the $ud$QM bag constant and the $a_{4}$ parameter. Specifically, set A corresponds to $B=20$ MeV/fm3 and $a_{4}=0.35$, set B to $B=35$ MeV/fm3 and $a_{4}=0.62$, and set C to $B=50$ MeV/fm3 and $a_{4}=0.88$.
@@ III Structure of Rotating stars
The structure of rotating neutron stars presents a significantly more complex problem to solve compared to their non-rotating counterparts. For non-rotating, spherically symmetric neutron stars, the combination of Einstein’s equation and the equilibrium condition yields the renowned Tolman-Oppenheimer-Volkoff Equations (TOV) Tolman (1939); Oppenheimer and Volkoff (1939).
@@
These equations must then be solved in all space, with flat space at infinity as boundary conditions. In this work we employ the code Astreus, which has been widely used to study rotating neutron stars Negreiros et al. (2017, 2014, 2013, 2012, 2010). This code employs the numerical method described in Komatsu et al. (1989) and have been thoroughly explored in Stergioulas and Friedman (1995).
We show in Fig. 2 the mass as a function of the central density corresponding to the three different EOSs for IHSs that we have considered.
Figure 2: Mass as a function of central density is depicted for three different Equations of State (EOS) of compact stars (IHSs) studied. For each EOS, two curves are presented, represented by solid and dashed lines, corresponding to the spherical sequence where stars have no rotation, and the Kepler sequence where each star rotates at its maximum (Kepler) frequency, respectively. Orange shaded regions represent unstable configurations
@@
The findings illustrated in Fig. 2 reveal that IHSs undergoing a transition from quark matter to hadronic matter at lower densities tend to exhibit smaller masses. Conversely, transitions occurring at higher densities result in larger masses.
@@
It is also worth noting the emergence of twin star configurations, especially for IHSs where the transition occurs at lower densities. Twin star configurations refer to scenarios where two stable stars from the same family have identical mass but differ in central densities (and radii), resulting in varying compactness Schramm et al. (2016); Dexheimer et al. (2015); Lyra et al. (2023). Typically, twin star setups involve the star with the smaller central density (less compact) being purely hadronic, while its more compact counterpart is hybrid. However, the case of inverted hybrid stars has the reversed situation, where the less compact twin is composed solely of quark matter, and the more compact twin contains both quark matter and hadronic matter, as explicitly shown in Zhang and Ren (2023) for non-rotating IHSs. Here our new finding reveals that, when rotation is introduced, an intriguing phenomenon emerges: the twin branch significantly expands. This is particularly noteworthy in the case of EOS B, where a relatively pronounced twin branch appears when rotation is factored in. This phenomenon may also appear in traditional hybrid stars. For instance, Ref. Tsaloukidis et al. (2023) showed that adding rotations can introduce twin stars from normal stars.
The emergence of twin-star configurations becomes more pronounced when examining the Mass-Radius diagram for these IHSs sequences. These findings are depicted in Figures 3, which display the gravitational mass as a function of the circumferential radius. The cyan-shaded region denotes the twin-star configurations. The insets in these figures illustrate the twin configurations for the corresponding non-rotating ones of the same EOS.
Figure 3: The mass as a function of circumferential radius for IHSs of EOS B (top panel) and EOS C (bottom panel) rotating at their Kepler frequencies. The blue shaded area denotes the zone where twin configurations occur. The insets illustrate the twin configurations of corresponding spherically symmetric (non-rotating) stars. Orange-shaded regions represent unstable configurations.
@@
Previous studies have also explored the possibility of twin stars in rotating configurations. Largani et al. (2022) investigated various models for rotating stars, including the novel aspect of considering finite temperatures. They observed the emergence of twin configurations; however, in their study, this phenomenon occurred with hot matter, rather than as a result of rotation, which is the focus of our present study. Another significant investigation into twin configurations in rotating stars is Bozzola et al. (2019). This work considered several microscopic models for hybrid stars, including those with a first-order phase transition. Notably, the models they explored exhibited prominent twin-star configurations in both rotating and nonrotating sequences. Despite their exemplary and comprehensive study of rotating hybrid stars, their findings do not indicate the emergence of twin stars solely due to rotation.
@@ IV Evolution Sequences
Given that a compact star may be born with a specific angular momentum, either after a supernova or a merger event, exploring potential rotational evolution paths as they decelerate is warranted. This involves examining the evolution of a sequence with constant baryon mass from a state of high rotation to the nonrotating limit. Such sequences can reveal the potential evolutionary trajectories of isolated stars that decelerate without mass loss or gain, thereby preserving a constant baryonic mass.
@@
Figure 4: Gravitational mass as a function of circumferential radius for IHSs of EOS B. Each dashed curve represents evolution paths for stars with the indicated baryon mass. The gray-shaded region depicts the forbidden region above the Kepler sequence. Similarly, orange shaded region represents unstable configurations.
It is crucial to note that while the stars depicted by the constant baryonic sequences in Fig. 4 are hydrostatically stable, not all may be stable against oscillations. The stability condition for a sequence of stars with constant baryonic mass follows the turning point theorem (see for instance Haensel et al. (2016)) and can be written as
@@
where $J$ is the angular momentum and $\epsilon_{c}$ the central energy density. We note that, strictly speaking, the condition given in eq. (9) should be used to determine secular stability, which may not necessarily coincide with dynamical stability Friedman et al. (1988). Besides the aforementioned condition, it is also necessary to consider the energy density gap in the EOS, which is evident in Fig. 1, resulting from the phase transition.
@@
Figure 6: Gravitational mass as a function of central density for the sequence of stars of EOS B. Black and red solid lines represent spherical and Kepler sequences, respectively. Dashed curves represent constant baryon mass sequences with the indicated baryonic mass, and orange-shaded regions represent forbidden configurations.
As observed in Fig. 6, there exists a small region of unstable configuration near the Kepler sequence. This suggests that stars may originate at lower frequencies (beyond the orange-shaded regions). Alternatively, if they are born with frequencies close to the Kepler limit, they will experience a sudden transition through the prohibited region. A more detailed analysis of the constant baryonic mass curves can provide a better understanding of this phenomenon, which can be facilitated with the aid of Fig. 7
Figure 7: Zoomed-in baryonic mass sequences shown in Fig. 6. Points A and B in each figure denote the last stable configuration before and after the forbidden region indicated by the orange shaded region.
@@
A similar result was outlined in reference Haensel et al. (2016). Similarly, references Chubarian et al. (2000); Mellinger et al. (2017) explored the possibility of phase transitions triggered by spin-down. However, in those cases, high-frequency stars in a purely hadronic phase underwent a phase transition to quark matter during spin-down. In contrast, our study finds that high-frequency stars are solely composed of quarks, with a phase transition to hadronic matter occurring during spin-down. The work of Ayvazyan et al. (2013) also concluded that heat would be released during spin-down triggered evolution; however, in their study, the phase transition was to a color-superconducting quark phase.
One must bear in mind, however, that not all of this energy is to be radiated away, as part of the energy will remain in the star in the form of heat. Even if only 1% of the energy is radiated it could give rise to very energetic phenomena such as gamma-ray bursts and fast-radio bursts Wang et al. (2024a), thus warranting further investigations on the dynamics of such phenomena.